OpenStudy (anonymous):

Solve 10x^4 -9x^2 +2=0 x=

OpenStudy (anonymous):

x = -sqrt(2/5) x = sqrt(2/5) x = -1/sqrt(2) x = 1/sqrt(2)

OpenStudy (anonymous):

so it should look like \[1/\sqrt{2}\]

OpenStudy (anonymous):

Yes, but simplify 1/sqrt{2} = sqrt{2}/2

OpenStudy (anonymous):

\[\sqrt{2/2}\]